# Spatial-frequency Dual-Domain Feature Fusion Network for Low-Light Remote Sensing Image Enhancement

### v1

@@ V-C3 Compared with transformer blocks
Figure 14: The visualization results compared with transformer blocks, ’One_block’ indicates that each block uses only one self-attention block and one FFN, while ’Stack_block’ signifies that multiple self-attention blocks and FFNs are stacked within each transformer block.
Initially, for fairness, we used only one Self-Attention block and one FFN operation per block, setting heads to 2. Although transformers are designed to capture long-range dependencies, shallow transformer blocks might struggle to learn rich feature representations and perform less effectively than frequency domain networks when dealing with large images. As shown in Table V, transformer blocks with only a single operation were not very effective and introduced a large number of parameters.
Subsequently, we followed the mainstream setup, stacking [2,4,6,8] self-attention blocks and FFN operations in each stage, with heads set to [2,2,4,4]. However, this setup increased the parameter count to nearly 27M, and the final enhancement effect was still slightly inferior to our method. In contrast, our frequency domain strategy only had 2.57M parameters but achieved the best results. Additionally, as shown in Fig. 14, the visualization demonstrates that our method achieves the best visual effects.
@@ VI Conclusion
Extensive quantitative and qualitative experiments demonstrate that the proposed method achieves outstanding performance in low-light remote sensing image enhancement, surpassing existing state-of-the-art methods on other datasets as well.

### v2

@@ V-C5 Compared with transformer blocks
Figure 17: Ablation Study on multistage loss. The left graph shows the loss decrease curves, and the right graph shows the PSNR increase curves. Each experiment was conducted for 1000 epochs.
Figure 18: The visualization results compared with transformer blocks, Single Transformer Block (STB) indicates that each block uses only one self-attention block and one FFN, while Multiple Transformer Block (MTB) signifies that multiple self-attention blocks and FFNs are stacked within each transformer block.
Initially, to ensure fairness, we used a single self-attention block and a single FFN operation in each block, setting num_head to 2. We refer to this as the Single Transformer Block (STB). Although Transformers are designed to capture long-term dependencies, shallower transformer blocks may struggle to learn rich feature representations, resulting in inferior performance compared to frequency domain networks when processing large images. As shown in Table VI, using only a single operation in transformer blocks leads to suboptimal results while introducing a significant number of parameters.
Subsequently, we followed the mainstream settings, stacking self-attention blocks and FFN operations in each stage as [2,4,6,8], with num_head set to [2,2,4,4], which we refer to as Multiple Transformer Block (MTB). However, this setup results in nearly 27M parameters, and the final enhancement effect is still slightly inferior to our method. Additionally, we measured the FLOPs and FPS on 1080p images for STB and our method (MTB could not run due to excessive computational requirements). In comparison, our frequency domain strategy has only 2.57M parameters, with 501 GFLOPs and 4 FPS, achieving the best results. Additionally, as shown in Fig. 18, the visualization demonstrates that our method achieves the best visual effects.
@@ VI Conclusion
Although our method demonstrates good performance in most cases, it still has certain limitations. The presence of numerous FFT and iFFT processes in the network lowers the model’s inference speed. Additionally, when there is a significant amount of noise in the environment, DFFN struggles to remove all the noise, which is a common challenge faced by current mainstream LLIE methods. In the future, we will focus on addressing the denoising problem in extremely low-light environments while improving the model’s inference speed, so that our method can be better applied to real-world scenarios.